OpenStudy (anonymous):

Write an equation for the line that passes through the points (0, -2) and (-4, 0). y= -2x- 4 y= -2x- 2 y= -1/2x - 2 y= -1/2x - 4 I believe it is y= -2x- 4, but I just want to make sure so I can see if I did something wrong.

jimthompson5910 (jim_thompson5910):

if you plugged x = 0 into y = -2x - 4, what would you get for y?

OpenStudy (anonymous):

It would be y= -2(0) - 4 I get, but I'm still confused.

jimthompson5910 (jim_thompson5910):

evaluate the right side

jimthompson5910 (jim_thompson5910):

do you know what I mean by that?

OpenStudy (anonymous):

Multiply -2(0)?

jimthompson5910 (jim_thompson5910):

to get what

OpenStudy (anonymous):

Y?

jimthompson5910 (jim_thompson5910):

-2 times 0 = ???

OpenStudy (anonymous):

0.

jimthompson5910 (jim_thompson5910):

so -2(0) - 4 turns into 0 - 4

jimthompson5910 (jim_thompson5910):

simplify 0-4 to get what number

OpenStudy (anonymous):

And you would simplify 0-4 by adding..I think, I'm confused again.

jimthompson5910 (jim_thompson5910):

no you just subtract

jimthompson5910 (jim_thompson5910):

0 - 4 = -4

jimthompson5910 (jim_thompson5910):

so if we started with y = 2x - 4, plugged in x = 0, we would get this y = -2x - 4 y = -2(0) - 4 y = 0 - 4 y = -4

jimthompson5910 (jim_thompson5910):

what does this mean?

OpenStudy (anonymous):

My answer was correct?

jimthompson5910 (jim_thompson5910):

no it means that the point (0,-4) is on the line but we want (0,-2) to be on the line

jimthompson5910 (jim_thompson5910):

see the difference?

OpenStudy (texaschic101):

would you use the slope formula first to find the slope ?

jimthompson5910 (jim_thompson5910):

yes that's one way to go about it texaschic101 but since we have answer choices, we can take advantage of that

OpenStudy (texaschic101):

ok...I understand

OpenStudy (anonymous):

No, I'm still confused a little.

jimthompson5910 (jim_thompson5910):

the idea is this: the two points must be on the line (both of them must be) so if you plugged each point one at a time, it should make the equation true (for both points)

OpenStudy (texaschic101):

would it be easier to show her the slope formula in case she has other questions that don't include answer choices ?

jimthompson5910 (jim_thompson5910):

y = 2x - 4 -2 = 2(0) - 4 ... plug in (0,-2) -2 = 0 - 4 -2 = -4 ... that's false, so (0,-2) is NOT on the line

jimthompson5910 (jim_thompson5910):

yes that's probably the best way to do it since multiple choice won't always be there

OpenStudy (anonymous):

Oh, I get that part now.

OpenStudy (texaschic101):

I am not trying to take over your explaining jim.....its all yours

jimthompson5910 (jim_thompson5910):

oh thanks, wasn't sure if you wanted to go or not lol

jimthompson5910 (jim_thompson5910):

anyways, the slope formula is this m = (y2 - y1)/(x2 - x1)

jimthompson5910 (jim_thompson5910):

the two points are: (0,-2) and (-4,0) so, (x1,y1) = (0,-2) ----> x1 = 0 and y1 = -2 (x2,y2) = (-4,0) ----> x2 = -4 and y1 = 0 we then plug all this into the slope formula

jimthompson5910 (jim_thompson5910):

m = (y2 - y1)/(x2 - x1) m = (0 - (-2))/(-4 - 0) m = (0 +2)/(-4 - 0) m = (2)/(-4) m = -1/2 so the slope is m = -1/2

OpenStudy (texaschic101):

Listen to him Kelli....he is good

jimthompson5910 (jim_thompson5910):

this means that the answer cannot be y = 2x - 4 since the slope for y = 2x - 4 is 2 (not -1/2) so we can effectively eliminate choices A and B

jimthompson5910 (jim_thompson5910):

with me so far?

OpenStudy (anonymous):

Yes, I'm with you so far, it's all just very confusing for me sometimes with all the numbers and etc.

jimthompson5910 (jim_thompson5910):

yeah it's a lot to keep track of

jimthompson5910 (jim_thompson5910):

but you'll get the hang of it

jimthompson5910 (jim_thompson5910):

we know that the slope is m = -1/2 one point is (0, -2), let's use this to find the y-intercept b y = mx + b y = (-1/2)x + b ... plug in the given slope -2 = (-1/2)(0) + b ... plug in (0,-2) -2 = 0 + b -2 = b b = -2 So the equation y = mx+b turns into y = -1/2x - 2

OpenStudy (anonymous):

I get it now, thank you.

jimthompson5910 (jim_thompson5910):

ok great, keep practicing it and you'll be a pro at it in no time